OpenStudy (anonymous):

Find the sum of the geometric series by using a formula? 1-4+16-64+256-1024+4096?

OpenStudy (tkhunny):

Rewrite: \(1 + 1\cdot (-4) + 1\cdot(-4)^{2} + 1\cdot(-4)^{3} + 1\cdot(-4)^{4} + 1\cdot(-4)^{5} + 1\cdot(-4)^{6}\)

OpenStudy (anonymous):

Do you know the formula for the sum of a finite geometric series?

OpenStudy (anonymous):

Yes it's Sn=a1 { (1+1)^n -1/1

OpenStudy (anonymous):

hmmm...may want to try that again, it looks a little off, unless I am misreading it.

OpenStudy (tkhunny):

I do, but I never use it. I just make it up on the fly. Let's see...\(\dfrac{1 - 1\cdot(-4)^{7}}{1-(-4)}\)

OpenStudy (anonymous):

\[Sn=a _{1} \left[ (1+1)^{n}-1 \right] / 1\]

OpenStudy (anonymous):

\[S _{n} = a _{1}\frac{ 1-r ^{n} }{ 1-r } \]

OpenStudy (anonymous):

Yes, yes, yes. sorry

OpenStudy (anonymous):

a1=1, n=7 and that's all we need.

OpenStudy (anonymous):

\[S _{7}=\frac{ 1-\left( -4 \right)7 }{ 1-(-4) }\] which becomes \[\frac{ 16385 }{ 5 }=3277\]

OpenStudy (anonymous):

ok, I see

OpenStudy (anonymous):

Good, coz all I see are commands...stupid pc is not enabling math...lol

OpenStudy (tkhunny):

Just learn where the formula comes from. FAR more useful than memorizing a formula!

OpenStudy (anonymous):

Agreed @tkhunny

OpenStudy (anonymous):

Thanks a lot guys!